OpenStudy (anonymous):

One of the roots of the equation 2x2 + 6x = 2k – 1 is twice the other. Find the value of k and the roots of the equation.

OpenStudy (anonymous):

can anybody help me?

OpenStudy (anonymous):

working on it. i think there is a snap way, but i don't see it yet. just grinding it out using quadratic formula

OpenStudy (anonymous):

but i am sure there is a simple way

OpenStudy (anonymous):

i've done it, but i want some answers to make sure that mine is correct

OpenStudy (anonymous):

use quadratic you get the two roots in terms of x, equate them solve for k so that 1 is double

OpenStudy (anonymous):

did you use the quadratic formula or did you use the sum and product of the roots? if you have an answer it is easy to check

OpenStudy (anonymous):

uhm, i have no answer for that to check it so.

OpenStudy (anonymous):

i use sum and product of the roots.

OpenStudy (anonymous):

it supposed to be 2x^2 , the question is a bit wrong

OpenStudy (anonymous):

i understand the question. you can grind it out by starting with \[2x^2+6x-2k+1=0\] then \[x=\frac{-6\pm\sqrt{36-8(-2k+1)}}{4}\] and work it that way. but i think it is silly. instead forget about the 2k-1 and just solve the problem with \[2x^2+6x+c=0\]

OpenStudy (anonymous):

so how to find the k and the other root?

OpenStudy (anonymous):

i must be doing something wrong, because if i try to solve it this way i get no solution

OpenStudy (anonymous):

are you sure it is \[2x^2+6x\] or might it be \[2x^2-6x\]?

OpenStudy (anonymous):

it is a plus

OpenStudy (anonymous):

ok let me try this \[2x^2+6x+c=0\] \[x=\frac{-6\pm\sqrt{36-8c}}{4}\] \[x=\frac{-6\pm2\sqrt{9-2c}}{2}\] \[x=\frac{-3\pm\sqrt{9-2c}}{2}\] so two roots are \[\frac{-3-\sqrt{9-2c}}{2}, \frac{-3+\sqrt{9-2c}}{2}\] and since one is twice the other you can write \[-\frac{3}{2}-\frac{\sqrt{9-2c}}{2}=-3+\sqrt{9-2c}\] or better still \[-3-\sqrt{9-2c}=-6+2\sqrt{9-2c}\] then \[3=3\sqrt{9-2c}\] \[\sqrt{9-2c}=1\]

OpenStudy (anonymous):

wow thats kind of hard way

OpenStudy (anonymous):

yes i agree, but i got \[c=\frac{9}{2}\] and let me check that it works

OpenStudy (anonymous):

nope it is wrong. damn

OpenStudy (anonymous):

how about the answer for k is 7/2 ?